#1) If tan of theta = (6/5) and cos of theta < 0, use the fundamental identities to evaluate the other five trigonometric functions of theta.
#2) use the fundamental identities to simplify csc squared if beta (1-cos squared of beta)
#3) factor and simplify. (sec "to the power of 4" x- tan "to the power of 4" x/ sec "squared" x + tan "squared" x)

Answer by stanbon(48558)

#1) If tan of theta = (6/5) and cos of theta < 0, use the fundamental identities to evaluate the other five trigonometric functions of theta.
---
If tan(theta) = 6/5 and cos(theta) < 0 then y = -6 and x = -5
-------------------------------
Solve for "r": r = sqrt(6^2+5^2) = sqrt(61)
---
sin(theta) = y/r = -6/sqrt(61)
cos(theta) = x/r = -5/sqrt(61)
tan(theta) = y/x = 6/5
-----------------
The csc, sec, and cot are the inverse of sin, cos, tan
================================================================

#2) use the fundamental identities to simplify csc squared of beta (1-cos squared of beta)
---
csc^2(1 - cos^2)
= csc^2 - (cos^2/sin^2)
= (1/sin^2) - (cos^2/sin^2)
= (1 - cos^2)/sin^2
= sin^2/sin^2
= 1
------------------------------------------------------
=================================================================
#3) factor and simplify. (sec "to the power of 4" x- tan "to the power of 4" x/ sec "squared" x + tan "squared" x)
-----------------------------------------
[sec^4(x) - tan^4(x)] / [sec^2(x) + tan^2(x)]
= [(sec^2+tan^2)(sec^2-tan^2)] / [sec^2 + tan^2]
= sec^2(x) - tan^2(x)
= 1
====================================================
Cheers,
Stan H.